# Thinking in Granularity: Dynamic Quantization for Image Super-Resolution by Intriguing Multi-Granularity Clues

### v1

@@ 3.3 Granular-DQ for SISR
Table 1: Quantitative comparison (FAB, PSNR (dB)/SSIM) with full precision models, PAMS, CADyQ, CABM, RefQSR and our method on Urban100, Test2K, Test4K for ×$\times 4$ SR.
@@
Table 2: Quantitative comparison (FAB, PSNR (dB)/SSIM) with full precision models, PAMS, CADyQ, CABM and our method on Urban100, Test2K, Test4K for ×$\times 2$ SR.
Specifically, we capture a generalized distribution statistic of the entropy for all LR patches on the training set. Given a patch $x$ of spatial size $h\times w$, its entropy is computed as
$\mathcal{H}(x)=-\sum^{255}_{\mu=0}\sum^{255}_{v=0}\mathcal{P}(\mu,v)\log{\mathcal{P}(\mu,v)},$ (8)
where $\mu$ and $v$ denote the current and neighbor pixel values respectively. $\mathcal{P}(\mu,v)=\mathcal{F}(\mu,v)/(hw)$ denotes the probability of $\mathcal{F}(\mu,v)$ manifesting within $x$ and $\mathcal{F}(\mu,v)$ signifies the frequency of occurrence of the tuple chuxia. In this way, we can get the entropy statistic across the overall training set, represented by $\mathbf{H}={\mathcal{H}_{1},\mathcal{H}_{2},...,\mathcal{H}_{M}}$ sorted in ascending order with $M$ patches, as shown in Fig. 5.
where $c_{n}$ denotes the adapted bit codes.
$t^{(j)}=t^{(j-1)}\cdot\gamma+{Norm(E)}\cdot(1-\gamma)$ (10)
where $Norm(\cdot)=\frac{\mathcal{H}_{t}-\mathcal{H}_{min}}{\mathcal{H}_{max}-\mathcal{H}_{min}}$, and $\mathcal{H}_{max}$ and $\mathcal{H}_{min}$ denotes the maximum and minimum entropy of all the patches in the current mini-batch at the $j$-th iteration. $\gamma$ represents the smoothing parameter of EMA, which is set to 0.9997. It should be noted that the LR samples remain consistent across epochs during training. Hence, our method only necessitates the E2B with ATC at the initial epoch, circumventing significant computational expenditure with iterations. Once the model is trained, as shown in Fig. 2, our method enables to fine-grained adapt the bit-widths of the patches based on calibrated thresholds from the large training set, yielding preferable bit codes $[c_{1},c_{2},...,c_{N}]$.
In summary, by combining GBC and E2B, our method ensures optimal bit allocation for each patch individually while dispensing with the consideration for layer sensitivity as previous methods [15, 39].
3.4 Loss Function
In previous SR quantization methods [15, 39, 25], the objective function is composed of L1 loss, knowledge distillation loss, and even bit regularization term to facilitate the bit adaption. In Granular-DQ, we only use $L_{1}$ loss to train all the models
$L_{1}=\left\|I_{HR}-I_{SR}\right\|_{1}$ (11)
4 Experiments
4.1 Experimental Settings
Baseline SR Models. The proposed Granular-DQ is applied directly to existing CNN-based SR models including SRResNet [24], EDSR [31], and IDN [18] as well as transformer-based models including SwinIR-light [30] and HAT-S [6]. Following CADyQ [15] and CABM [39], we implement quantization on the weights and feature maps within the high-level feature extraction part, which is the focal point for the majority of computationally intensive operations.
In Granular-DQ, the first step for bit allocation by GBC designates 4/6/8-bit as the candidate bits to quantize the patches. Subsequently, the second step by E2B adapts the patches allocated with 8 bits in GBC are further adapted using 4/5/8-bit as the candidates for fine-grained bit-width adjustment. The initial entropy thresholds, denoted as $t_{1}$ and $t_{2}$, are set to 0.5 and 0.9 respectively and then gradually calibrated according to the entropy statistic on the training set, for all models. In this work, we employ QuantSR [34] for all the quantization candidates and uniformly apply 8-bit linear quantization for weights.
Datasets and Metrics. In our experiments, all the models are trained on DIV2K [1] dataset which contains 800 training samples for ×$\times 2$ and ×$\times 4$ SR. We evaluate the model and compare it with existing methods on three benchmarks: Urban100 [17], Test2K and Test4K [22] derived from DIV8K dataset [12] by bicubic downsampling. We quantitatively measure the SR performance using two metrics: peak signal-to-noise ratio (PSNR) and the structural similarity index (SSIM) for reconstruction accuracy. Besides, we also compute the feature average bit-width (FAB) which represents the average bit-width across all features within the test dataset to measure the quantization efficiency.
Implementation details. For the transformer-based models, the linear layers of the MLPs in both SwinIR-light [30] and HAT-S [6] are all quantized using the QuantSR scheme [34]. Notably, for SwinIR-light and HAT-S, the attention blocks are computed with full precision due to severe quantization errors. During training, we randomly crop each LR RGB image into a $48\times 48$ patch with a batch size of 16 and $64\times 64$ with a total batch size of 16 for CNN- and transformer-based baselines respectively. All the models are trained for 300K iterations on NVIDIA RTX 4090 GPUs with Pytorch. The learning rate is set to $2\times 10^{-4}$ and is halved after 250K iterations. During testing, the input image is split into $96\times 96$ LR patches.
4.2 Comparing with the State-of-the-Art
Quantitative Comparison. Tab. 1 reports the quantitative results on benchmarks. The proposed Granular-DQ is compared with original full-precision models, PAMS [28], CADyQ [15], CABM [39], AdaBM [14], and RefQSR [25]. One can see that Granular-DQ demonstrates the minimum performance sacrifice relative to the full-precision SRResNet and EDSR models while attaining the lowest FAB against other methods on all benchmarks. For IDN, Granular-DQ even exceeds its full-precision model by about 0.2dB on Urban100 and Test4K datasets, whereas other methods show lower PSNR and SSIM improvements with obviously higher FAB. We also demonstrate the comparison on scaling factor ×$\times 2$ in Tab. 2. Obviously, Granular-DQ demonstrates competitive trade-offs in terms of FAB and PSNR/SSIM compared to other quantization methods across all CNN models. Moreover, when implementing these methods on transformer-based baselines, it can be observed that Granular-DQ significantly outperforms other methods in terms of reconstruction accuracy and quantization efficiency. The results validate the superior effectiveness and generalization ability of Granular-DQ.
Method FAB Params (K) BitOPs (G)
(↓ Ratio) (↓ Ratio)
EDSR 32.00 1518K (0.0%) 527.0T (0.0%)
PAMS 8.00 631K (↓ 58.4%) 101.9T (↓ 80.7%)
CADyQ 6.09 489K (↓ 67.8%) 82.6T (↓ 84.3%)
CABM 5.80 486K (↓ 68.0%) 82.4T (↓ 84.4%)
Ours 4.97 486K (↓ 68.0%) 73.6T (↓ 86.0%)
Qualitative Comparison. Figure 6 shows the qualitative results on the Urban100 dataset. As one can see, Granular-DQ produces SR images with sharper edges and clearer details, sometimes even better than the original unquantized IDN. By comparison, despite the lower PSNR and more FAB consumption, existing methods also suffer from obvious blurs and misleading textures. More qualitative results see in the Fig. 7 and Fig. 8.
Complexity Analysis. To further investigate the complexity of our method for quantizing SR models, we calculate the number of operations weighted by the bit-widths (BitOPs) [41] as the metric and compare it with existing methods. As shown in Tab. 3, Granular-DQ leads to significant computational complexity reduction of the baseline model, which decreases the BitOPs from 527.0T to 73.6T and sustains a competitive FAB. Coupled with the decrease in the model parameters to 68.0% (486K) of the full-precision model, the results demonstrate that Granular-DQ can ensure optimal trade-off between reconstruction accuracy and quantization efficiency.
4.3 Ablation Study
Effects of Individual Components. We study the effects of the proposed components including GBC, E2B, and ATC in Tab. 4, where the results are evaluated on the Urban100 dataset. We can see that quantization with only GBC leads to a performance drop. Based on GBC, when we introduce E2B to conduct fine-grained bit-width adaption, the resultant quantizer can enhance the reconstruction accuracy and a small improvement in efficiency. Moreover, E2B and ATC in conjunction effectively reduce the FAB by a considerable margin (over 0.5 FAB) with almost the same PSNR/SSIM.
GBC E2B ATC Urban100
FAB PSNR SSIM
8.00 26.01 0.783
5.86 25.97 0.782
5.51 26.02 0.784
4.97 26.01 0.784
$b^{*}$ Set14 Urban100
FAB PSNR SSIM FAB PSNR SSIM
$[4,5,6]$ 5.29 28.52 0.780 4.85 25.98 0.783
$[4,5,7]$ 5.50 28.54 0.780 4.98 25.99 0.782
$[4,6,7]$ 5.79 28.55 0.781 5.22 25.97 0.783
$[4,6,8]$ 5.64 28.57 0.781 5.38 26.01 0.784
$[4,7,8]$ 5.64 28.55 0.780 5.64 25.99 0.783
$[4,5,8]$ 5.54 28.58 0.781 4.97 26.01 0.784
$t_{1}$ $t_{2}$ Set14 Urban100
FAB PSNR SSIM FAB PSNR SSIM
0.4 0.7 5.86 28.53 0.780 5.28 25.96 0.783
0.4 0.8 5.86 28.50 0.779 5.04 25.99 0.783
0.4 0.9 5.54 28.56 0.780 4.99 25.98 0.782
0.5 0.7 5.86 28.53 0.780 5.25 25.97 0.784
0.5 0.8 5.82 28.57 0.780 5.02 26.00 0.782
0.5 0.9 5.54 28.58 0.781 4.97 26.01 0.784
Table 6: Ablation study on the impact of the thresholds in ATC with EDSR baseline.
$t^{*}$ $b^{*}$ Set14 Urban100
FAB PSNR SSIM FAB PSNR SSIM
$[0.5]$ $[4,8]$ 6.57 28.57 0.780 5.75 25.97 0.781
$[0.5,0.9]$ $[4,5,8]$ 5.54 28.58 0.781 4.97 26.01 0.784
$[0.4,0.6,0.9]$ $[4,5,6,8]$ 6.07 28.58 0.779 5.41 25.93 0.781
$[0.4,0.6,0.9]$ $[4,5,7,8]$ 6.21 28.54 0.780 5.61 25.93 0.782
Table 7: Ablation study on the influence of a different number of thresholds (quantile, denoted by $t^{*}$) and corresponding bit configuration (denoted by $b^{*}$) in E2B with EDSR.
Influence of the Candidate Bits in E2B. We conduct experiments to investigate the influence of the bit configuration in E2B. For 3 candidate bits, we set the lowest bit-width as 4 and randomly change the other two, resulting in 6 variants. As reported in Tab. 5, the configuration of [4, 5, 6] performs worst on both Set14 and Urban100 with relatively lower FAB. Surprisingly, although we allocate higher bit-width to patches ([4, 7, 8]), the model incurs the most FAB but acquires negligible performance gains. By comparison, the model with [4, 5, 8] achieves the best trade-off on the two datasets, which is selected as our final configuration.
Impact of the Threshold $t$ in ATC. In this work, we set two thresholds $t_{1}$ and $t_{2}$ in ATC, which divide the entropy of input patches into 3 subintervals and then map them to the bit codes ([4, 5, 8] in Tab. 5), whitch facilitates the bit-width adjustment in E2B. As reported in Tab. 6, according to the results on Set14 and Urban100, we can empirically set the combination of [$t_{1}=0.5$, $t_{2}=0.9$] as it achieves the best balance in quantization.
Impact of the Threshold Number in E2B. We further experimentally investigated the effect of different numbers of thresholds in E2B and their corresponding candidate bit configuration. Firstly, we assume that there is only one quantile for all the input patches, which means the entropy statistic $\mathbf{H}$ is divided into two subintervals. As shown in Tab. 7, when we adjust the bit-widths of patches using 4/8bit, the model performs worst on both Set14 and Urban100 datasets. Similarly, when we incorporate three thresholds of $t$ with $[0.4,0.6,0.9]$ to divide $\mathbf{H}$ into four subintervals, it can be seen that whether using the bit configurations of $[4,5,6,8]$ or $[4,5,7,8]$, the model cannot obtain satisfied quantization efficiency. In contrast, the model with two thresholds $[0.5,0.9]$ and corresponding candidate bit-widths of $[4,5,8]$ achieved the best trade-off on both datasets, making it our final choice.
Influence of Different Quantization Patterns. To investigate the compatibility of our method on different quantization patterns, we conduct experiments by combining Granular-DQ with PAMS [28] and QuantSR [34], where the results on Urban100 are reported in Tab. 8. Notably, different from existing dynamic methods [15, 39], our Granular-DQ does not require the pre-trained models of PAMS or QuantSR. We can see that Granular-DQ+PAMS gets 0.07dB PSNR gains with 0.4 FAB reduction for EDSR compared to CADyQ+PAMS. When applying the QuantSR scheme on Granular-DQ, the model can achieve the best trade-off between FAB and PSNR/SSIM for both EDSR and IDN models, where even the latter surpasses the original model by 0.26dB in PSNR.
Methods Urban100
FAB↓ PSNR↑ SSIM↑
EDSR 32.00 26.03 0.784
PAMS 8.00 26.01 0.784
CADyQ+PAMS 6.09 25.94 0.782
Granular-DQ+PAMS 5.69 25.95 0.782
Granular-DQ+QuantSR 4.97 26.01 0.784
IDN 32.00 25.42 0.763
PAMS 8.00 25.56 0.768
CADyQ+PAMS 5.78 25.65 0.771
Granular-DQ+PAMS 4.73 25.62 0.770
Granular-DQ+QuantSR 4.18 25.68 0.772
Table 8: Investigation of the compatibility of our Granular-DQ with different quantization patterns. We observe the ×$\times 4$ SR results on Urban100 based on EDSR and IDN.
5 Limitation
@@
6 Conclusion

### v2

@@ Introduction
Table 1: Quantitative comparison (FAB, PSNR (dB)/SSIM) with full precision models, PAMS, CADyQ, CABM, RefQSR and our method on Urban100, Test2K, Test4K for ×$\times 4$ SR. ×$\times 2$ SR results are provided in the supplementary material.
Specifically, we capture a generalized distribution statistic of the entropy for all LR patches on the training set. We first discretize the total $N$ pixels within a patch into multiple bin intervals $B$ based on the pixel values, which can estimate the probability distribution of pixels smoothly. Then entropy is computed as
$\mathcal{H}=-\sum_{i=1}^{N}\mathcal{P}(x_{i})log(\mathcal{P}(x_{i})).\vspace{-0.1cm}$ (8)
We use Gaussian-weighted kernel to assign different importance to the pixels in a patch with the formulation of ${\sum_{i=1}^{N}\sum_{j=1}^{B}exp(-\frac{(r_{i})^{2}}{2\sigma^{2}})+\epsilon}$, where $r_{i}$ denotes the residual between the pixel value of the $i$-th pixel $x_{i}$ and the segment values for bin intervals. Thus, one can obtain its kernel density $\mathcal{P}(x_{i})$ by $\frac{\sum_{j=1}^{B}exp(-\frac{(r_{i})^{2}}{2\sigma^{2}})}{\sum_{i=1}^{N}\sum_{j=1}^{B}exp(-\frac{(r_{i})^{2}}{2\sigma^{2}})+\epsilon}$. In this way, we can get the entropy statistic across the overall training set, represented by $\mathbf{H}={\mathcal{H}_{1},\mathcal{H}_{2},...,\mathcal{H}_{M}}$ sorted in ascending order with $M$ patches, as shown in Figure 4.
where $c_{n}$ denotes the adapted bit codes.
$t^{(j)}=t^{(j-1)}\cdot\gamma+{Norm(E)}\cdot(1-\gamma),$ (10)
where $Norm(\cdot)=\frac{\mathcal{H}_{t}-\mathcal{H}_{min}}{\mathcal{H}_{max}-\mathcal{H}_{min}}$, and $\mathcal{H}_{max}$ and $\mathcal{H}_{min}$ denotes the maximum and minimum entropy of all the patches in the current mini-batch at the $j$-th iteration. $\gamma$ represents the smoothing parameter of EMA, which is set to 0.9997. It should be noted that the LR samples remain consistent across epochs during training. Hence, our method only necessitates the E2B with ATC at the initial epoch, circumventing significant computational expenditure with iterations. Once the model is trained, as shown in Figure 2, our method enables to fine-grained adapt the bit-widths of the patches based on calibrated thresholds from the large training set, yielding preferable bit codes $[c_{1},c_{2},...,c_{N}]$.
In summary, by combining GBC and E2B, our method ensures optimal bit allocation for each patch individually while dispensing with the consideration for layer sensitivity as previous methods (Hong et al. 2022a; Tian et al. 2023).
Loss Function
In previous SR quantization methods (Hong et al. 2022a; Tian et al. 2023; Lee, Yoo, and Jung 2024), the objective function is composed of L1 loss, knowledge distillation loss, and even bit regularization term to facilitate the bit adaption. In Granular-DQ, we only use $L_{1}$ loss to train all the models
$L_{1}=\left\|I_{HR}-I_{SR}\right\|_{1}$ (11)
Experiments
Experimental Settings
Baseline SR Models. The proposed Granular-DQ is applied directly to existing CNN-based SR models including SRResNet (Ledig et al. 2017), EDSR (Lim et al. 2017), and IDN (2018) as well as transformer-based models including SwinIR-light (Liang et al. 2021) and HAT-S (Chen et al. 2023). Following CADyQ (Hong et al. 2022a) and CABM (Tian et al. 2023), we implement quantization on the weights and feature maps within the high-level feature extraction part, which is the focal point for the majority of computationally intensive operations. Notably, for SwinIR-light and HAT-S, the attention blocks are computed with full precision due to severe quantization errors, where more details are provided in the supplementary material.
In Granular-DQ, the first step for bit allocation by GBC designates 4/6/8-bit as the candidate bits to quantize the patches. Subsequently, the second step by E2B adapts the patches allocated with 8 bits in GBC are further adapted using 4/5/8-bit as the candidates for fine-grained bit-width adjustment. The initial entropy thresholds, denoted as $t_{1}$ and $t_{2}$, are set to 0.5 and 0.9 respectively and then gradually calibrated according to the entropy statistic on the training set, for all models. In this work, we employ QuantSR (Qin et al. 2024) for all the quantization candidates and uniformly apply 8-bit linear quantization for weights.
Datasets and Metrics. In our experiments, all the models are trained on DIV2K (Agustsson and Timofte 2017) dataset which contains 800 training samples for ×$\times 2$ and ×$\times 4$ SR. We evaluate the model and compare it with existing methods on three benchmarks: Urban100 (Huang, Singh, and Ahuja 2015), Test2K and Test4K (Kong et al. 2021) derived from DIV8K dataset (Gu et al. 2019) by bicubic downsampling. We quantitatively measure the SR performance using two metrics: peak signal-to-noise ratio (PSNR) and the structural similarity index (SSIM) for reconstruction accuracy. Besides, we also compute the feature average bit-width (FAB) which represents the average bit-width across all features within the test dataset to measure the quantization efficiency.
Implementation details. During training, we randomly crop each LR RGB image into a $48\times 48$ patch with a batch size of 16. All the models are trained for 300K iterations on NVIDIA RTX 4090 GPUs with Pytorch. The learning rate is set to $2\times 10^{-4}$ and is halved after 250K iterations. During testing, the input image is split into $96\times 96$ LR patches.
Comparing with the State-of-the-Art
Quantitative Comparison. Table  1 reports the quantitative results on benchmarks. The proposed Granular-DQ is compared with original full-precision models, PAMS (Li et al. 2020), CADyQ, CABM, AdaBM (Hong and Lee 2024), and RefQSR (2024). One can see that Granular-DQ demonstrates the minimum performance sacrifice relative to the full-precision SRResNet and EDSR models while attaining the lowest FAB against other methods on all benchmarks. For IDN, Granular-DQ even exceeds its full-precision model by about 0.2dB on Urban100 and Test4K datasets, whereas other methods show lower PSNR and SSIM improvements with obviously higher FAB. Moreover, when implementing these methods on transformer-based baselines, it can be observed that Granular-DQ significantly outperforms other methods in terms of reconstruction accuracy and quantization efficiency. The results validate the superior effectiveness and generalization ability of Granular-DQ.
Qualitative Comparison. Figure 5 shows the qualitative results on the Urban100 dataset. As one can see, Granular-DQ produces SR images with sharper edges and clearer details, sometimes even better than the original unquantized IDN. By comparison, despite the lower PSNR and more FAB consumption, existing methods also suffer from obvious blurs and misleading textures.
Method FAB Params (K) BitOPs (G)
(↓ Ratio) (↓ Ratio)
EDSR 32.00 1518K (0.0%) 527.0T (0.0%)
PAMS 8.00 631K (↓ 58.4%) 101.9T (↓ 80.7%)
CADyQ 6.09 489K (↓ 67.8%) 82.6T (↓ 84.3%)
CABM 5.80 486K (↓ 68.0%) 82.4T (↓ 84.4%)
Ours 4.97 486K (↓ 68.0%) 73.6T (↓ 86.0%)
Complexity Analysis. To further investigate the complexity of our method for quantizing SR models, we calculate the number of operations weighted by the bit-widths (BitOPs) (Van Baalen et al. 2020) as the metric and compare it with existing methods. As shown in Table 2, Granular-DQ leads to significant computational complexity reduction of the baseline model, which decreases the BitOPs from 527.0T to 73.6T and sustains a competitive FAB. Coupled with the decrease in the model parameters to 68.0% (486K) of the full-precision model, the results demonstrate that Granular-DQ can ensure optimal trade-off between reconstruction accuracy and quantization efficiency.
Ablation Study
Effects of Individual Components. We study the effects of the proposed components including GBC, E2B, and ATC in Table 3, where the results are evaluated on the Urban100 dataset. We can see that quantization with only GBC leads to a performance drop. Based on GBC, when we introduce E2B to conduct fine-grained bit-width adaption, the resultant quantizer can enhance the reconstruction accuracy and a small improvement in efficiency. Moreover, E2B and ATC in conjunction effectively reduce the FAB by a considerable margin (over 0.5 FAB) with almost the same PSNR/SSIM.
GBC E2B ATC Urban100
FAB PSNR SSIM
8.00 26.01 0.783
5.86 25.97 0.782
5.51 26.02 0.784
4.97 26.01 0.784
$b^{*}$ Set14 Urban100
FAB PSNR SSIM FAB PSNR SSIM
$[4,5,6]$ 5.29 28.52 0.780 4.85 25.98 0.783
$[4,5,7]$ 5.50 28.54 0.780 4.98 25.99 0.782
$[4,6,7]$ 5.79 28.55 0.781 5.22 25.97 0.783
$[4,6,8]$ 5.64 28.57 0.781 5.38 26.01 0.784
$[4,7,8]$ 5.64 28.55 0.780 5.64 25.99 0.783
$[4,5,8]$ 5.54 28.58 0.781 4.97 26.01 0.784
Influence of the Candidate Bits in E2B. We conduct experiments to investigate the influence of the bit configuration in E2B. For 3 candidate bits, we set the lowest bit-width as 4 and randomly change the other two, resulting in 6 variants. As reported in Table 4, the configuration of [4, 5, 6] performs worst on both Set14 and Urban100 with relatively lower FAB. Surprisingly, although we allocate higher bit-width to patches ([4, 7, 8]), the model incurs the most FAB but acquires negligible performance gains. By comparison, the model with [4, 5, 8] achieves the best trade-off on the two datasets, which is selected as our final configuration. More ablations are provided in the supplementary material.
Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62472137, 62072151, 62302141, 62331003, 62206003), Anhui Provincial Natural Science Fund for the Distinguished Young Scholars (2008085J30), Open Foundation of Yunnan Key Laboratory of Software Engineering (2023SE103), CCF-Baidu Open Fund (CCF-BAIDU202321), CAAI-Huawei MindSpore Open Fund (CAAIXSJLJJ-2022-057A) and the Fundamental Research Funds for the Central Universities (JZ2024HGTB0255).
Supplementary Material
Thinking in Granularity: Dynamic Quantization for Image Super-Resolution by Intriguing Multi- Granularity Clues
Mingshen Wang1, Zhao Zhang1,2*, Feng Li1*, Ke Xu3, Kang Miao1, Meng Wang1
1Hefei University of Technology
2Yunnan Key Laboratory of Software Engineering
3Anhui University
Appendix
The supplementary material mainly includes the following contents:
  • The motivation of Granular-DQ.
  • Ablation studies involve the thresholds (quantile $t$) number in E2B, and the combination with different weight quantization patterns, i.e. PAMS (Li et al. 2020) and QuantSR (Qin et al. 2024).
  • More implementation details on the transformer-based baseline models including SwinIR-light (Liang et al. 2021) and HAT-S (Chen et al. 2023).
  • More experimental results consist of ×$\times 2$ SR performance and additional qualitative visualization.
  • Limitations in Granular-DQ.
Motivation
Recent advances (Tu et al. 2023; Hong et al. 2022a; Tian et al. 2023; Lee, Yoo, and Jung 2024; Hong and Lee 2024) have demonstrated the benefits of considering the quantization sensitivity of layers and image contents in SR quantization. Taking CADyQ (Hong et al. 2022a) for example, it applies a trainable bit selector to determine the proper bit-width and quantization level for each layer and a given local image patch based on the feature gradient magnitude. In our analysis, we compute the average bit-width and the quantization error measured by MSE between the reconstructions of the quantized model (via CADyQ) and the original high-precision model (EDSR) on the Test2K dataset. Figure S1 (a) reveals that the majority of patches fall within a 6-bit to 8-bit range, accompanied by a relatively elevated MSE. Furthermore, we present t-SNE maps for various quantized layers and the final layer in Figure S1 (c)-(d). Firstly, it is evident that the distribution of different layers quantized by CADyQ is markedly more scattered than that of the original model, as depicted in Figure S1 (c). Secondly, on the final layer, the features from the CADyQ-quantized model exhibit a distinct vertical pattern, which is notably at odds with the structure of the original model’s feature points (Figure S1 (d)). In our investigation, CABM actually exhibited similar findings, although it fine-tunes the CADyQ model based on edge scores. These results indicate that: 1) Simply relying on image edge information is suboptimal for the trade-off between quantization efficiency and error; 2) The bit allocation for each layer in response to varying patches can introduce disturbances to the inter-layer relations within original models leading to disparities in the representations.
Figure S1: Analysis of the quantization efficiency, quantization error, and feature distribution in t-SNE on CADyQ and our Granular-DQ. (a) and (b) illustrate the quantization efficiency v.s. quantization error trade-off; (c) and (d) visualize the feature distribution of two resultant models and compare with the corresponding original one (Float32: EDSR).
Based on the above analysis, this work aims to design a dynamic quantization approach for diverse image contents while maintaining the representation ability of the original model. To this end, we rethink the image characteristics related to image quality from the granularity and information density. As we know, the fine-granularity representations reveal the texture complexity of local regions, while coarse ones express structural semantics of the overall scene. Besides, according to Shannon’s Second Theorem (Shannon 1948), the entropy statistic reflects the average information density and the complexity of pixel distributions given patches, which is directly correlated to the image quality. Therefore, we propose Granular-DQ, a markedly different method that fully explores the granularity and entropy statistic of images to quantization adaption. Granular-DQ contains two sequential steps: 1) granularity-aware bit allocation for all the patches and 2) entropy-based fine-grained bit-width adaption for the patches less quantized by 1). In this way, we can see that the bit-width allocation by Graular-DQ is sparser than CADyQ, where a majority of patches are lower than 5bit with only a few patches at high bit-width (Figure S1 (b)). Moreover, the feature distribution of the layers quantized by our method is closer to that of the original model (Figure S1(c)-(d)). These validate that our Granular-DQ enables low-bit and layer-invariant quantization.
$t_{1}$ $t_{2}$ Set14 Urban100
FAB PSNR SSIM FAB PSNR SSIM
0.4 0.7 5.86 28.53 0.780 5.28 25.96 0.783
0.4 0.8 5.86 28.50 0.779 5.04 25.99 0.783
0.4 0.9 5.54 28.56 0.780 4.99 25.98 0.782
0.5 0.7 5.86 28.53 0.780 5.25 25.97 0.784
0.5 0.8 5.82 28.57 0.780 5.02 26.00 0.782
0.5 0.9 5.54 28.58 0.781 4.97 26.01 0.784
Table S1: Ablation study on the impact of the thresholds in ATC with EDSR baseline.
$t^{*}$ $b^{*}$ Set14 Urban100
FAB PSNR SSIM FAB PSNR SSIM
$[0.5]$ $[4,8]$ 6.57 28.57 0.780 5.75 25.97 0.781
$[0.5,0.9]$ $[4,5,8]$ 5.54 28.58 0.781 4.97 26.01 0.784
$[0.4,0.6,0.9]$ $[4,5,6,8]$ 6.07 28.58 0.779 5.41 25.93 0.781
$[0.4,0.6,0.9]$ $[4,5,7,8]$ 6.21 28.54 0.780 5.61 25.93 0.782
Table S2: Ablation study on the influence of a different number of thresholds (quantile, denoted by $t^{*}$) and corresponding bit configuration (denoted by $b^{*}$) in E2B with EDSR.
Ablation Study
Impact of the Threshold $t$ in ATC. In this work, we set two thresholds $t_{1}$ and $t_{2}$ in ATC, which divide the entropy of input patches into 3 subintervals and then map them to the bit codes ([4, 5, 8] in Table S1), whitch facilitates the bit-width adjustment in E2B. As in Table S1, according to the results on Set14 and Urban100, we can empirically set the combination of [$t_{1}=0.5$, $t_{2}=0.9$] as it achieves the best balance in quantization.
Impact of the Threshold Number in E2B. We further experimentally investigated the effect of different numbers of thresholds in E2B and their corresponding candidate bit configuration. Firstly, we assume that there is only one quantile for all the input patches, which means the entropy statistic $\mathbf{H}$ is divided into two subintervals. As shown in Table S2, when we adjust the bit-widths of patches using 4/8bit, the model performs worst on both Set14 and Urban100 datasets. Similarly, when we incorporate three thresholds of $t$ with $[0.4,0.6,0.9]$ to divide $\mathbf{H}$ into four subintervals, it can be seen that whether using the bit configurations of $[4,5,6,8]$ or $[4,5,7,8]$, the model cannot obtain satisfied quantization efficiency. In contrast, the model with two thresholds $[0.5,0.9]$ and corresponding candidate bit-widths of $[4,5,8]$ achieved the best trade-off on both datasets, making it our final choice.
Influence of Different Quantization Patterns. To investigate the compatibility of our method on different quantization patterns, we conduct experiments by combining Granular-DQ with PAMS (Li et al. 2020) and QuantSR (Qin et al. 2024), where the results on Urban100 are reported in Table S3. Notably, different from existing dynamic methods (Hong et al. 2022a; Tian et al. 2023), our Granular-DQ does not require the pre-trained models of PAMS or QuantSR. We can see that Granular-DQ+PAMS gets 0.07dB PSNR gains with 0.4 FAB reduction for EDSR compared to CADyQ+PAMS. When applying the QuantSR scheme on Granular-DQ, the model can achieve the best trade-off between FAB and PSNR/SSIM for both EDSR and IDN models, where even the latter surpasses the original model by 0.26dB in PSNR.
Methods Urban100
FAB↓ PSNR↑ SSIM↑
EDSR 32.00 26.03 0.784
PAMS 8.00 26.01 0.784
CADyQ+PAMS 6.09 25.94 0.782
Granular-DQ+PAMS 5.69 25.95 0.782
Granular-DQ+QuantSR 4.97 26.01 0.784
IDN 32.00 25.42 0.763
PAMS 8.00 25.56 0.768
CADyQ+PAMS 5.78 25.65 0.771
Granular-DQ+PAMS 4.73 25.62 0.770
Granular-DQ+QuantSR 4.18 25.68 0.772
Table S3: Investigation of the compatibility of our Granular-DQ with different quantization patterns. We observe the ×$\times 4$ SR results on Urban100 based on EDSR and IDN.
Implementation Details on Transformer-based Baselines
For the transformer-based models, the linear layers of the MLPs in both SwinIR-light (Liang et al. 2021) and HAT-S (Chen et al. 2023) are all quantized using the QuantSR scheme (Qin et al. 2024). Surprisingly, despite all our efforts, we still encountered the gradient explosion issue when implementing the quantization scheme for CADyQ (Hong et al. 2022a) and CABM (Tian et al. 2023) in HAT-S. As a result, these two retained full precision for channel attention in the experiments. During the training phase, we randomly cropped the LR image into 64 $\times$ 64 with a total batch size of 16 for all scale factors, following the settings of the original model. The learning rate was initially set to $2\times 10^{-4}$ and halved after 250K iterations.
@@ Acknowledgments
Table S4: Quantitative comparison (FAB, PSNR (dB)/SSIM) with full precision models, PAMS, CADyQ, CABM and our method on Urban100, Test2K, Test4K for ×$\times 2$.
Comparison with the State-of-the-Art
Quantitative Comparison for $\times$2 SR. We further conduct experiments for ×$\times 2$ SR, where the quantitative results are illustrated in Table S4. Obviously, Granular-DQ demonstrates competitive trade-offs in terms of FAB and PSNR/SSIM compared to other quantization methods across all CNN models. Additionally, for SwinIR-light and HAT-S, Granular-DQ also achieves remarkably superior reconstruction accuracy to full precision than others while maintaining the lowest FAB.
More Qualitative Comparison. In Figure S2 and S3, we provide more ×$\times 4$ SR visual results produced by recent state-of-the-art methods and our Granular-DQ. Regardless of whether the models are CNN-based or transformer-based, Granular-DQ consistently achieves superior reconstruction details at the lowest FAB compared to other quantization methods in most instances. In each case, minimal discrepancies can be observed between Granular-DQ and its corresponding full-precision model. These findings further validate that Granular-DQ ensures an optimal trade-off between reconstruction accuracy and quantization efficiency.
Limitation
@@
Figure S2: More visual comparison ($\times$4) on Urban100 (×$\times 4$) for different methods.
Figure S3: More visual comparison ($\times$4) on Test2K (×$\times 4$) for different methods.